OpenStudy (help_people):

Sam is observing the velocity of a car at different times. After two hours, the velocity of the car is 50 km/h. After six hours, the velocity of the car is 54 km/h. Part A: Write an equation in two variables in the standard form that can be used to describe the velocity of the car at different times. Show your work and define the variables used. (5 points) Part B: How can you graph the equation obtained in Part A for the first seven hours? (5 points)

OpenStudy (help_people):

@welshfella

OpenStudy (help_people):

@Nnesha @pooja195

OpenStudy (help_people):

thank you so much @dumbcow for helping me :)

OpenStudy (dumbcow):

get equation of a line from two points the 2 points given are: x = hours y = mph (2,50) (6,54) m =slope = (54-50)/(6-2) = 4/4 = 1 y = mx + b ----> y = 1x +b b = y - x = 50 -2 = 48 equation of line is: y = x + 48

OpenStudy (help_people):

Thank you who about for Part B @dumbcow

OpenStudy (help_people):

*how

OpenStudy (help_people):

@ganeshie8 may you finish part b please?

OpenStudy (help_people):

or @Data_LG2

OpenStudy (anonymous):

dumbcow already gave the equation. All you have to do is find is the velocity of the car after 1hr, 2hr, 3hr, 4hr, 5hr, 6hr and 7hr. The x variable is your time and y is your velocity. Plug the values of x to find the values of y. After you find those values, you can easily graph the data.

OpenStudy (help_people):

for part b ?

OpenStudy (anonymous):

yep

OpenStudy (help_people):

wait so i have to graph every sing one i am so confused?

OpenStudy (help_people):

@Data_LG2

OpenStudy (help_people):

like y=1+48 and y=2+48, y=7+48

OpenStudy (anonymous):

yes :)

OpenStudy (help_people):

so how would i put this in a paragraph.

OpenStudy (help_people):

I can first start by saying to plug in values for the first seven hours. I will change the x value and replace it with 1, 2, 3, 4, 5, 6, and 7?

OpenStudy (help_people):

Then I could show the equations maybe is thats what it is axing?